OpenStudy (anonymous):

The length of a rectangle is 4 times its width and the perimeter is 150 feet. What is the width of the rectangle? A. 60 feet B. 15 feet C. 75 feet D. 30 feet

OpenStudy (anonymous):

@timaashorty

OpenStudy (timaashorty):

L (2) + W (2)= perimeter So we know that Length = 4 * W Since you have answer choices I would suggest you to use them A. 60 * 4 = 240 ✘ B. 15*4 = 60 ; So 15(2) + 60(2) = 150 ✓ B is correct (:

OpenStudy (anonymous):

thank you

OpenStudy (timaashorty):

welcome